OpenStudy (anonymous):

Hey guys.... can some one explain me what a newton is, and how do I find net force step by step...thank you.

OpenStudy (fretje):

The Newton is the Force a body of 1kg undergoes when it is accelerated by 1m/s2 . So unit of Newton is \[N = kg.\frac{m}{s^2}\]

OpenStudy (anonymous):

whats confusing me is M/^2 I mean, thats really fast... but thanks any way.

OpenStudy (fretje):

what is confusing you?

OpenStudy (anonymous):

metre per secon squared

OpenStudy (anonymous):

so after you multiply the kilogram, what do you do?

OpenStudy (anonymous):

I'm trying to find out How much net force is required to accelerate a 1000 kg car at 5.00 m/s2

OpenStudy (fretje):

metre per secon squared means every second, the speed goes up with one meter per second, so with one Newton, the kg goes after first second 1m/s, after 2s 2m/s, after 3s 3m/s and so on. It is comparable to about one tenth of the gravitational acceleration which is 9,8 m/s

OpenStudy (anonymous):

oooh, so if you had 2 metres it would be 2 seconds....ok..

OpenStudy (fretje):

Sir Isaac Newton had a rule to calculate any force: F (the force) = m.a (is the mass in kg, times the acceleration in m/s2)

OpenStudy (anonymous):

aaaahhh, ok, I finally get it, thank you!

OpenStudy (fretje):

I'm trying to find out How much net force is required to accelerate a 1000 kg car at 5.00 m/s2 So F=m.a = 1000kg.5.00m/s2 = 5000 Newton